OpenStudy (anonymous):

Canned fruits, vegetables, and soups have labels on the curved surface of the metal container. What is the surface area of the label on this can? That is, what is the can's lateral area? A cylinder with radius eight centimeters and height eleven centimeters is shown. eighty-eight square centimeters four hundred one point nine two square centimeters five hundred fifty-two point six four square centimeters nine hundred fifty-four point six four square centimeters

OpenStudy (anonymous):

Well the diameter of a circle is 2 *pi * radius If you multiply that by the height of the cylinder that should give you the answer!

OpenStudy (anonymous):

so 69.08?

OpenStudy (anonymous):

Should be a bit more than that,\[2 \times r \times \pi \times h\]We know that r=6 and that h=11

OpenStudy (anonymous):

Oh Okay! So 414.48?

OpenStudy (anonymous):

Yeah that would be right, but I misread the question and r = 8

OpenStudy (anonymous):

Ok thank you! i got it now.